how do I solve 2cosxcscx-4cosx-cscx+2=0
thank you!

Answer by lwsshak3(11628)

how do I solve 2cosxcscx-4cosx-cscx+2=0
**
0 < x < π/2
2cosxcscx-4cosx-cscx+2=0
cscx(2cosx-1)-2(2cosx-1)=0
(2cosx-1)(cscx-2)=0
(2cosx-1)=0
cosx=1/2
x=π/3
or
cscx-2=0
cscx=2
sinx=1/2
x=π/6